### v2

# Demystifying Misconceptions in Social Bots Research

Demystifying Misconceptions in Social Bots Research
Stefano Cresci,1,∗ Kai-Cheng Yang,2 Angelo Spognardi,3 Roberto Di Pietro,4
Filippo Menczer,5 Marinella Petrocchi1
1IIT-CNR, Italy
2Northeastern University, USA
3“La Sapienza” University of Rome, Italy
4KAUST, Saudi Arabia
5Indiana University, USA
Research on social bots aims at advancing knowledge and providing solutions to one of the most debated forms of online manipulation. Yet, social bot research is plagued by widespread biases, hyped results, and misconceptions that set the stage for ambiguities, unrealistic expectations, and seemingly irreconcilable findings. Overcoming such issues is instrumental towards ensuring reliable solutions and reaffirming the validity of the scientific method. In this contribution, we review some recent results in social bots research, highlighting and revising factual errors as well as methodological and conceptual biases. More importantly, we demystify common misconceptions, addressing fundamental points on how social bots research is discussed. Our analysis surfaces the need to discuss research about online disinformation and manipulation in a rigorous, unbiased, and responsible way. This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
Corresponding author: stefano.cresci@iit.cnr.it
Introduction
Human decision-making processes depend on the availability of high quality information. In addition, a healthy society requires a shared understanding of issues and values to foster trust and cooperation. Misinformation and disinformation erode trust and emphasize divisions. Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach consensus and take effective collective action. The repercussions could be so severe as to delay responses to a deadly pandemic (Cinelli et al., 2020; Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
The science of misinformation seeks solutions to these problems (Lazer et al., 2018). However, this paramount endeavor can itself incur in the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023). For example, scientific articles and publishers engage in a fierce competition for attention, similarly to what is being experienced by news outlets. As a consequence, sensationalist claims and hyped results can sometimes be used as shortcuts to publication and other scientific accomplishments (West and Bergstrom, 2021). Then, confirmation bias is making its way into science in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021). And again, over-generalizations of scientific results and poor understanding of methodological and conceptual limitations give rise to multiple misconceptions about misinformation (Altay et al., 2023). These and other problems currently undermine the efficacy and credibility of our research efforts. Therefore, for the science of misinformation to benefit our society, we must first solve the problems within the science itself.
The present study concerns one of the many forms of online disinformation: malicious social bots. Social bots are accounts on an online platform that are—at least partially—automated (Ferrara et al., 2016; Cresci, 2020). Thanks to their ease of use and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content. Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb misinformation and disinformation (Shao et al., 2018). For this reason, social bots have attracted considerable scholarly and media attention (Allem and Ferrara, 2018; Assenmacher et al., 2020; González-Bailón and De Domenico, 2021; Chen et al., 2021). One of the latest such episodes was the public dispute on the real number of bots on Twitter that took place before Elon Musk’s acquisition of the platform (Varol, 2023). However, despite many years of research, the science of social bots is replete with the same problems that also plague the science of misinformation. These originate from uncorrected factual errors, from the presence of biases in how scientific results are cited and discussed, and from a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
In our present analyses, we focus on some recently published articles about social bots. A thorough analysis of their methods, assumptions, results, and conclusions reveals several pitfalls ranging from factual inaccuracies and methodological flaws to more subtle incidents such as misrepresentation of the existing literature and conceptual errors. Our discussion links these issues to common misconceptions in social bots and misinformation research. Our critical analysis thus serves a twofold goal. On the one hand, we highlight and revise factual errors, methodological biases, and conceptual issues. On the other hand, demystifying common misconceptions allows us to address fundamental issues on how science is produced and discussed (West and Bergstrom, 2021). Overcoming such issues is instrumental towards ensuring the credibility of science.
Methodological issues
Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Ferrara et al., 2016; Cresci, 2020). In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots—on various online platforms. The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts with the goal of detecting subtle differences between authentic users and automated entities. Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
Information leakage
Information leakage in machine learning refers to situations where a model is exposed to confidential information from the training data that is later exploited to make unfair inferences (Chakrabarti et al., 2008). Feature leakage is a specific type of information leakage where the model exploits machine learning features that are not available at prediction time. This issue typically results in the trained model exhibiting unrealistically good performance due to the exploitation of “forbidden” information. In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance. For example, each botnet typically exhibits some peculiar characteristics that surface as a result of how the bots were created, their goals, or any other shared characteristic between them (Zhang et al., 2016; Mazza et al., 2022). These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Barabasi, 2005; Cresci et al., 2020). However, most of the time, such peculiar characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets. Therefore, if a bot detector exploits such characteristics at training time for classifying instances of known bots, it might learn spurious correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
Another way in which feature leakage occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise). Again, the correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading. Consequently, a detector exploiting these features would achieve excellent performance on known data but would likely exhibit poor capacity to generalize to new data with different characteristics.
The inadvertent or deliberate inclusion of information leakage in a bot detector’s training process introduces a potential avenue for its authors to artificially elevate its performance, enabling strategic competition with established state-of-the-art models. Such situation can lead to the proposal of ostensibly superior models, which however represent a regression rather than advancement in addressing the bot detection problem. In addition, the same models could also be exploited to unduly undermine the performance of other models. This was recently done by Gallwitz and Kreil (2022) and Hays et al. (2023), who trained simple but unrealistic bot detectors using spurious signals, such as whether an account was verified by Twitter, to exaggerate the limitations of more general state-of-the-art models.
Cherry-picking
Another challenge in bot detection arises from the multitude of disparate detectors and benchmark datasets (Cresci, 2020). This creates an environment where authors introducing new detectors can selectively engage in favorable comparisons. The issue is exacerbated by the difficulty of delineating a clear state-of-the-art in bot detection, as the sheer volume of existing models makes it arduous to discern the most effective ones. The lingering uncertainty around the performance of even established bot detectors (Rauchfleisch and Kaiser, 2020) makes it possible to cherry-pick competitors and evaluation scenarios, allowing proponents of a novel detector to demonstrate its superiority only against a small subset of detectors and datasets, conveniently omitting those against which it may not perform as well.
To make matters worse, despite the existence of high-quality research published in prestigious venues (Shao et al., 2018; Stella et al., 2018; González-Bailón and De Domenico, 2021), the heightened interest in social bots has attracted a wide spectrum of contributions, including some lower-quality works. For example, Gallwitz and Kreil (2022) selected non-representative samples (prominent public accounts, such as members of the US Congress and Nobel Prize Laureates) to challenge the accuracy of a widely-used bot detection tool, inflating its false-positive rate. Such misleading claims underscore the fallacy of cherry-picking evaluation data. Since bot detection methods, whether based on machine learning or on human annotation, have inherent error margins, accuracy results can be manipulated by carefully choosing test examples. This less rigorous body of work, when added to the literature, creates a reference that, while possibly lacking in quality, further facilitates the selective cherry-picking of competitors.
The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that benefit from information leakage or that overfit to specific evaluation datasets. [^1] We refer to overfitting when a model depends too much on irrelevant features of the training data, with the result that it performs well on the training data but generalizes poorly to unseen data (Bramer, 2007). State-of-the-art models are designed for greater generality, rendering direct comparisons with these simpler yet contextually specialized models inherently unfair. In a recent example of this kind, overfitted classifiers with as few as four features are compared to more complex and general state-of-the-art models (Hays et al., 2023).
The issue of cherry-picking extends beyond the unfair comparisons between bot detectors. A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings. Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023) and to falsely claim that bot detectors have not been systematically evaluated (Gallwitz and Kreil, 2022). These claims can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020). On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—to give a false impression of prior research supporting their claims. Such selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art in bot detection. By self-citing unvetted claims, authors create an illusion of authority. And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed to the same or comparable insights. This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of researchers who have previously advanced the field.
Addressing the cherry-picking issue necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature. Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts. This discrepancy surfaces a crucial tension in maintaining the quality and rigor of peer review processes within rapidly evolving research landscapes (Van Noorden, 2023). The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of questionable quality.
Straw-man methodology
The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation as if it faithfully represented the original research. For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on misrepresentation of the literature. As part of their criticism of supervised methods, they falsely claim that other researchers do not share or publish training and evaluation datasets. They also make demonstrably false claims about other researchers manipulating results through the use of unreported training data or threshold hacking. Gallwitz and Kreil criticize supervised bot detectors as invalid on the basis of these serious claims of scientific misconduct, yet do not provide any evidence in support of their allegation. Another methodological straw-man fallacy is the claim that Gallwitz and Kreil failed to reproduce an experiment by Keller and Klinger (2019). Gallwitz and Kreil applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in the bot detector.
A different manifestation of the straw-man fallacy is the claim that bot detection is exclusively a supervised machine learning task. On the one hand, this assumption could be motivated by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of available supervised bot detectors (Cresci, 2020). On the other hand, such a view overlooks a similarly robust literature highlighting that supervised detectors suffer from lack of generalizability and transferability, as well as other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). A variety of unsupervised approaches to bot detection have recently emerged as possible solutions to mitigate these problems (Cresci, 2020). Evaluating the generalization capabilities of bot detectors is precisely the goal of Hays et al. (2023). However, they chose to do so by adopting a supervised learning methodology that has been proven to suffer from critical issues when it comes to generalizability and transferability. Another issue with their methodology is that it does not fully support their conclusions. Specifically, Hays et al. leverage their results to criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods. They do not make any distinction between the different types of bot detectors proposed to date—supervised ones based on the analysis of individual accounts vs. unsupervised ones based on the analysis of groups of accounts (Cresci, 2020). This is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable. However, the analysis only evaluates supervised detectors, which have known generalization issues (Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). The unsurprising—and biased—result is that the considered detectors fail to generalize, which is used to criticize the whole field.
Data biases
In machine learning and data science tasks, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009). As mentioned above, numerous datasets of bot and human accounts have been published over the years. On the one hand, this made it easier to thoroughly test the performance of new bot detectors. On the other hand, it introduced the risk of cherry-picking favorable datasets, potentially distorting reported performances and introducing bias. Furthermore, the temporal heterogeneity stemming from the availability of datasets spanning varying time periods—ranging from recent to decade-old—can pose additional challenges when detectors are trained on outdated data, undermining the relevance of their performance in contemporary settings. When considering the representativeness of published benchmark datasets with respect to the bots that currently inhabit online platforms, the readers should be aware that the accounts included in a dataset are likely a few years older than the publication date of the dataset itself. Moreover, some datasets are built by combining or extending previously published resources. For example, the pan-2019 dataset (Rangel and Rosso, 2019) is largely based on other datasets published between 2011 and 2017. Hence, the publication date of a dataset should be considered as a generous upper bound of the recency of the accounts therein. Even under this relaxed assumption, in consideration of the rapid pace of change of the landscape of online harms and of the availability of more comprehensive and more recent datasets (Feng et al., 2021), we may run the risk of capturing datasets that are no longer representative of the actual state of the platforms. This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021).
Some recent literature investigates the characteristics of accounts that are over ten years old. Obsolete datasets can be used to criticize other people’s detection methods or to promote one’s own method as better than the state of the art. A glaring example is the work of Hays et al. (2023), where out of eleven analyzed datasets, the most recent was published in 2020 and the least recent in 2011.
Annotations can be another source of data bias. For example, so-called cyborg accounts (Ferrara et al., 2016) are automated in part via software, but closely managed and monitored by humans who can step in as needed to avoid detection. By labeling such accounts as human on the basis that they can reply to a query and assert that they are not bots, one can make it appear that a bot detection method has a higher false-positive rate than it actually does. To avoid this kind of bias, account labeling should be performed by multiple annotators following a shared and published rubric. Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) that accounts exhibit should be favored in place of binary labels.
Some work in the literature fails to follow this methodology, only to accuse other researchers of bias. Gallwitz and Kreil (2022) demonstrate this behavior in several cases. Without disclosing any annotation rubric, they change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative. In one case, criticizing Yang et al. (2019), they label accounts that automatically post news headlines, such as those associated with major newswire agencies, as human. Their criticism of Keller and Klinger (2019) goes further: first, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger; second, they manually inspected only 0.04% of this different set; third, they ignored suspended accounts; and, fourth, they labeled as human accounts those that cross-posted tweets through software apps. Their conclusion was that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word.» Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al. (2020): suspended accounts are again assumed to be human to boost the alleged false-positive rate, then accounts that are clearly automated are still labeled as classification errors by changing the standard from “bot” to “malicious” or “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions is a dangerous trend toward scientific manipulation.
(a) Simplistic and naïve model of social bots and human-operated accounts. Here, all bots are alike, they are evenly distributed, and we have complete knowledge on their numbers and characteristics. According to this model, collecting unbiased and comprehensive bot datasets is possible.
(b) Realistic model of social bots and human-operated accounts, and of the related knowledge gaps. Here, social bots are organized in botnets, each with peculiar characteristics (color-coded). Among the missing information is the number of bots in known botnets, as well as the number, size, and characteristics of all unknown botnets.
Figure 1: The ecology of online accounts: social bots and humans. Differences between naïve and realistic models, and their implications for data collection.
Conceptual issues
Beyond methodological challenges, the landscape of social bots research is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues. These nuanced challenges, if left unaddressed, have the potential to exert an even more insidious influence on the development of the field and its reception.
Failure to account for context
In discussing the limitations that currently hinder progress in bot detection, some recent research has pointed the finger at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023). This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in datasets acquired with rigorous practices, contrary to the literature so far. More specifically, this criticism of the existing data collection practices subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient. Unbiased datasets are needed so as to avoid that peculiar bot and human characteristics leak into the data allowing simple models to achieve good performance on the collected data, while showing poor performance in real scenarios. Furthermore, comprehensive datasets are needed so that expressive models can learn to generalize to all types of existing social bots. Unfortunately, the assumption of the existence of an unbiased and comprehensive dataset is fallacious.
Regarding bias, many bot datasets were found to contain biases, in that the accounts therein featured some peculiar characteristics (Hays et al., 2023). We note that this is largely unsurprising since such peculiar characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a single entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016). Given that all bots belonging to the same botnet are created and/or operated by the same entity, and pursue the same common goal, they tend to share some similarities. When compared to other accounts, the botnet thus appears to have some peculiar characteristics, as highlighted in Figure 1(b). Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather should be traced back to the very nature of the social bot phenomenon.
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance. Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020). we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. In addition to the many challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021; Torres-Lugo et al., 2022). Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^3] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question cannot currently be answered.
Common misconceptions in social bots research
Part of the social bot literature claims that the bot detection task has been effectively solved. Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications. This optimistic view positions bot detection research as a success story for machine learning, highlighting the progress made in achieving practical and deployable solutions in the field.
We disagree with this interpretation of the state of social bot research. On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. We are well aware of this body of work, having actively contributed to its development for many years. Below we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media. [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020; Pacheco et al., 2021). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
Misconception 1 Social bot detection is a solved task.
The above evidence shows that despite significant efforts devoted for a prolonged time (Cresci, 2020), bot detection is nowhere near to being a solved problem—quite the contrary.
In framing the bot detection task as a seemingly straightforward problem, some authors also hint at the possibility of easily adapting existing detectors to overcome their current limitations and keep pace with the evolution of more human-like bots (Hays et al., 2023). This example introduces another misconception about bot detection:
Misconception 2 Bot detection performance can be improved easily.
For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Cresci et al., 2021). Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology. Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
Below we discuss some other misconceptions that undermine our understanding of social bots and of the related literature:
Misconception 3 All social bots are similar.
Misconception 4 Each bot detector can detect all types of bots.
A common assumption in the scientific literature, as well as among the general public, is that if a given bot detector has achieved good detection performance against some bots, then that detector should be able to replicate that performance on any bot detection task. This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022) and only considering the known botnets. Consider, for example, the differences between the bots used to boost the popularity of certain public figures—so-called fake followers (Cresci et al., 2015)—and those that manipulate trending topics—astroturfers and spammers (Abokhodair et al., 2015). Or the differences between bots involved in political manipulation (Shao et al., 2018; Caldarelli et al., 2020) versus those aimed at fooling automated trading systems (Tardelli et al., 2022; Bello et al., 2023). In other words, not all bots are the same. Some scholars addressed this heterogeneity by designing detectors that aim for generality and broad applicability, such as Botometer, especially in its latest releases (Sayyadiharikandeh et al., 2020; Yang et al., 2020). Others have developed detectors specifically designed to detect certain types of bots. The latter trade generalizability and portability for detection accuracy. Their performance depends heavily on the characteristics of the bots they are designed to detect. For example, a detector designed to detect time-synchronized retweeting actions (Mazza et al., 2019) would likely be useless at detecting mass-following bots (Cresci et al., 2015). However, this should not be regarded as a limitation of the bot detector—let alone one deliberately unstated by its developers in order to pass it off as a good product—but rather as an inappropriate use of the detector itself. Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized detectors. In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot or not. Therefore, in general, any detector, specialized or otherwise, has variable performance and its detection capabilities depend on the characteristics of the accounts it is applied to. In conclusion, no single bot detector is capable of detecting all types of bots.
Many articles proposing a new bot detector, or investigating bot interference, motivate the work on the basis of the role of social bots in spreading disinformation (Shao et al., 2018). Similar arguments are made for the role of bots in exacerbating social conflicts (Stella et al., 2018). In other words, part of the social bot literature is based on the assumption that bots are responsible for many of the problems that affect our online platforms and the information ecosystem as a whole. This brings to the fore another misconception:
Misconception 5 Social bots are mainly responsible for the spread of disinformation.
An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020; Pierri et al., 2023). Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
The previous misconceptions might induce readers to think that social bot research has led to flawed—if not outright useless—results. Such a claim has been made by Hays et al. (2023). Similarly, Gallwitz and Kreil (2022) not only question the validity of existing bot detection methods—based on the fallacies discussed earlier—but go even further in arguing that social bots are not an issue. These claims lead to our last, but by no means least important, misconception:
Misconception 6 All social bots research is flawed and bot detection results are useless.
We see at least two strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the findings. These cases are just some of the success stories of social bot research. If researchers had not developed bot detection techniques, we would not have been able to identify pockets of anomalous accounts engaged in sophisticated manipulations. Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none. Second, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Ratkiewicz et al., 2011; Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
Challenges in the post-API era
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. Previously, studies primarily concentrated on Twitter/X, likely because its free API facilitated data acquisition. However, this accessibility ceased in 2023 when Twitter/X terminated free data access for researchers. [^7] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^8] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence in real time. It significantly impedes the collection of new bot samples necessary for studying their characteristics and training novel machine learning classifiers. Even if researchers can develop new classifiers, deploying them at scale is challenging, exposing social media users to potential manipulation. Bot operators, on the other hand, remain largely unaffected. Leveraging burner and virtual cellphones, they can manage bot accounts across platforms directly, bypassing API restrictions. The advent of AI-powered social bots, which evade current bot detection models (Yang and Menczer, 2023; Yang et al., 2024), further emphasizes the importance of data availability in support of social bot research.
Despite these setbacks, Europe’s Digital Service Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^9] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question. Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community.
Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
A call for moral responsibility
Our analysis of the drawbacks and limitations of some recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars—such as those between proponents and opponents of social bots research.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019). Indeed, we believe that all the literature that propagates the misconceptions demystified in this paper provides misleading contexts in some way. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.
Making tangible progress in the science of misinformation requires that we address many conceptual, practical, and ethical challenges. Doing so involves embracing the intrinsic complexity of the phenomenon, considering multiple viewpoints, and providing nuance rather than naivety. Sweeping statements such as “the creators of bot datasets are responsible for the failure of the field,” “all social bots research is flawed or useless,” or even “this new bot detector has flawless performance” do not contribute to achieving these goals.
In conclusion, should we fail to respect our moral obligation, we would produce biased and unreliable research, further worsening the problems that currently undermine the credibility of our field (West and Bergstrom, 2021; Altay et al., 2023). In a 2019 piece published in Science, Derek Ruths compared contrasting findings about social bots, commenting that «research on misinformation has come to resemble the very thing it studies» (Ruths, 2019). It is our moral responsibility to reverse this fatal course. This can only be achieved via responsible research that fosters nuanced, unbiased, and balanced viewpoints, and by a review process that adheres to the same principles. This article aims to make a contribution in this direction by debunking common fallacious arguments adopted by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; by the European Union – Horizon 2020 Program under the scheme “INFRAIA-01-2018-2019 – Integrating Activities for Advanced Communities”, Grant Agreement n. 871042, “SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics”; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.